OpenStudy (anonymous):

cos(x)-sin(x)=sqrt(2)sin(x) prove that: cos(x)+sin(x)=sqrt(2)cos(x)

OpenStudy (anonymous):

is this the question \[\cos(x)-\sin(x)= \sqrt2 sinx\] prove \[cosx+sinx = \sqrt2 cosx\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Any specific way you need to prove?

OpenStudy (anonymous):

No

OpenStudy (usukidoll):

I'm confused . That could be \[\sqrt{2}sinx\]or \[\sqrt{2sinx}\] :/

OpenStudy (anonymous):

\[\sqrt{2} sinx\]

OpenStudy (usukidoll):

alright... there's got to be an identity we can use for this

OpenStudy (anonymous):

don't no ... i'm also confused